OpenStudy (anonymous):

(5.6-2a+1.4)/1.2=4.5 please help

OpenStudy (paki):

(5.6-2a+1.4)/1.2=4.5 5.6-2a+1.4=4.5*1.2... solve for "a" now...

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

thank you

OpenStudy (paki):

pleasure